# OmniAudio: Generating Spatial Audio from 360-Degree Video

### v1

@@ Abstract
Traditional video-to-audio generation techniques primarily focus on field-of-view (FoV) video and non-spatial audio, often missing the spatial cues necessary for accurately representing sound sources in 3D environments. To address this limitation, we introduce a novel task, 360V2SA, to generate spatial audio from 360-degree videos, specifically producing First-order Ambisonics (FOA) audio - a standard format for representing 3D spatial audio that captures sound directionality and enables realistic 3D audio reproduction. We first create Sphere360, a novel dataset tailored for this task that is curated from real-world data. We also design an efficient semi-automated pipeline for collecting and cleaning paired video-audio data. To generate spatial audio from 360-degree video, we propose a novel framework OmniAudio, which leverages self-supervised pre-training using both spatial audio data (in FOA format) and large-scale non-spatial data. Furthermore, OmniAudio features a dual-branch framework that utilizes both panoramic and FoV video inputs to capture comprehensive local and global information from 360-degree videos. Experimental results demonstrate that OmniAudio achieves state-of-the-art performance across both objective and subjective metrics on Sphere360. Code and datasets will be released at https://github.com/liuhuadai/OmniAudio. The demo page is available at https://OmniAudio-360V2SA.github.io.
360-degree video-to-spatial audio, spatial audio generation, conditional flow matching
@@ 1 Introduction
{subcaptionblock}
0.49 Refer to caption {subcaptionblock}0.49 Refer to caption
Figure 1: Comparison of panoramic video and perspective video.
@@ 2 Related Work
Video-to-Audio Generation. This direction (Zhang et al., 2024; Wang et al., 2024a; Xu et al., 2024c; Tian et al., 2024) focuses on synthesizing audio that aligns seamlessly with the visual content of a video clip. Some approaches, such as SpecVQGAN (Iashin & Rahtu, 2021), FoleyGen (Mei et al., 2024), and V-AURA (Viertola et al., 2024), leverage autoregressive techniques to produce audio from silent video. Meanwhile, other methods (Luo et al., 2024; Cheng et al., 2024b; Xu et al., 2024b) adopt diffusion models (Song et al., 2020; Ho et al., 2020; Xing et al., 2024) or flow matching (Lipman et al., 2022; Vyas et al., 2023; Wang et al., 2024b) generative models. Diff-Foley (Luo et al., 2024) employs an audio-visual contrastive feature and latent diffusion to predict spectrogram latent. Similarly, MovieGen Audio (Polyak et al., 2024) uses flow matching conditioned on multi-modal inputs, including videos and texts. Despite these advancements, a significant gap remains in the research focused on generating spatial audio from 360-degree video, an essential component for crafting genuinely immersive experiences. We introduce OmniAudio, which, to the best of our knowledge, represents the first approach in spatial audio generation from 360-degree video.
@@
Spatial Audio Generation. Existing spatial audio  [^1] Binaural audio captures sound using two microphones to simulate human ear positioning for a 3D effect when using headphones, while spatial audio involves a broader range of techniques that create 3D sound in dynamic environments, including with head tracking or multi-speaker setups. generation methods (Xu et al., 2021; Leng et al., 2022; Liu et al., 2023b; Gao & Grauman, 2019b; Kushwaha et al., 2024; Anonymous, 2024) predominantly focus on producing binaural (Sun et al., 2024; Yoshida et al., 2023; Garg et al., 2021) or First-order Ambisonics (FOA) audio (Kushwaha et al., 2024; Dagli et al., 2024) from fixed perspective inputs, such as mono audio, visual features, text, and source locations. For example, 2.5D Visual Sound (Gao & Grauman, 2019b) transforms monoaural audio into binaural sound using convolutional neural networks, while Diff-SAGe (Kushwaha et al., 2024) and ViSAGe (Anonymous, 2024) generate FOA audio by incorporating sound categories, source locations, and camera parameters. However, these approaches typically assume a static camera perspective, limiting their applicability in dynamic environments where the field of view can change in real-time. To overcome these limitations, we propose a framework for generating FOA audio from 360-degree video, which inherently supports dynamic and panoramic visual inputs. Our method leverages self-supervised flow matching pre-training and dual-branch video design to effectively synthesize spatial audio that adapts to the comprehensive and evolving visual context captured by 360-degree videos. This advancement enables more immersive and flexible audio-visual experiences in applications requiring real-time spatial audio adjustments.
@@
Flow Matching. Flow matching is used as the backbone for audio generation due to its superior generation performance. This framework (Lipman et al., 2022) trains a model to learn a vector field, enabling the generation of a desired probability path from noise to data. Unlike score-based models such as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), FM offers more stable and robust training and enhanced performance. This generative model has proven effective in audio generation, as seen in applications like Audiobox (Vyas et al., 2023), AudioLCM (Liu et al., 2024a), and FlashAudio (Liu et al., 2024b). Moreover, SpeechFlow (Liu et al., 2023a) has integrated self-supervised pre-training with FM to improve speech processing tasks, demonstrating its adaptability. In our work, we adapt FM for the novel task of spatial audio generation from 360-degree videos. By adopting a coarse-to-fine self-supervised pre-training strategy, our approach first learns general audio patterns from large-scale non-spatial audio data and then fine-tunes the model to capture spatial-specific characteristics. This enables the generation of high-quality First-order Ambisonics (FOA) audio that complements the immersive visual context of 360-degree videos.
@@ 4.1 Preliminaries
Conditional Flow Matching. Conditional flow matching is an effective generative technique applied across image, audio, and video domains (Lipman et al., 2022). It employs a time-dependent velocity vector field parameterized by a neural network, denoted as $v_{\theta}(t,C,x)$, where $t$ represents time, $C$ is the conditioning input (e.g., video or text), and $x$ is a data point in $\mathbb{R}^{d}$.
@@ 4.2 Spatial Audio Representation
To address this gap and effectively encode FOA audio, we propose several modifications based on the Stable Audio framework: (1) We initialize our four-channel VAE with weights from a pre-trained stereo VAE to leverage existing non-spatial audio knowledge. (2) We eliminate the Mid-Side Short Time Fourier Transform (MS-STFT), which is specifically designed for stereo reconstruction, and adapt the system to transform the left/right components into FOA components W, X, Y, and Z, each with an equal loss weight of 1/4. For detailed information on VAE, please refer to Appendix D.
@@ 4.3 Dual-Branch Video Representation
We propose a dual-branch architecture to effectively model spatial-temporal dynamics in 360-degree videos. Given an input video sequence $V_{360}\in\mathbb{R}^{T\times H\times W\times C}$, which is in equirectangular projection format  [^3] The Equirectangular Projection (ERP) is a popular method for projecting 360-degree videos, commonly used by platforms like YouTube. It maps a spherical view onto a 2D image using latitude and longitude coordinates, with a 2:1 aspect ratio (360° horizontal, 180° vertical). with a height-to-width ratio of 1:2, we first pad it to a square format (1:1) and then resize it to match the input requirements of the image encoder. Here, $T$ is the number of frames, $H$ is the frame height, $W$ is the frame width, and $C$ is the number of color channels. We extract the FOV video, denoted as $V_{\text{fov}}\in\mathbb{R}^{T\times H^{\prime}\times W^{\prime}\times C}$, from the 360-degree video $V_{360}$ to capture local information. Our framework employs two complementary branches leveraging a frozen pre-trained MetaCLIP-Huge image encoder (Xu et al., 2024a):
@@ 5 Experiments
Table 2: Performance comparison between OmniAudio and the baselines on the Sphere360-Bench test set and YT360 test set. We use objective metrics computing FD, KL divergence, $\Delta_{abs}\theta$, $\Delta_{abs}\phi$, and $\Delta_{\text{Angular}}$ between estimated DoA and ground truth, as well as subjective metrics including MOS for spatial audio quality (MOS-SQ) and video-audio alignment faithfulness (MOS-AF). We report the mean and standard deviation for MOS-SQ and MOS-AF. +AS denotes adding an audio spatialization component. For metrics with a downward arrow ($\downarrow$), lower values represent better performance, while for metrics with an upward arrow ($\uparrow$), higher values indicate better quality.
@@
GT YT360-Test / / / / / 85.38$\pm$0.95 87.85$\pm$1.21
Diff-Foley + AS YT360-Test 361.65 2.22 / / / 67.21$\pm$0.95 70.34$\pm$1.76
MMAudio + AS YT360-Test 190.40 1.71 / / / 73.25$\pm$1.05 76.77$\pm$1.23
ViSAGe (FOV) YT360-Test 199.09 1.86 2.21 0.88 1.99 71.82$\pm$1.98 72.17$\pm$1.47
ViSAGe (360) YT360-Test 225.52 1.95 2.18 0.86 1.98 72.45$\pm$1.64 72.96$\pm$1.39
OmniAudio YT360-Test 92.57 1.64 1.27 0.53 1.27 80.37$\pm$0.91 83.49$\pm$1.01
GT Sphere360-Bench / / / / / 88.41$\pm$0.79 90.12$\pm$1.08
Diff-Foley + AS Sphere360-Bench 331.05 3.56 / / / 69.87$\pm$0.84 71.12$\pm$1.36
MMAudio + AS Sphere360-Bench 271.15 2.39 / / / 75.34$\pm$0.99 77.56$\pm$1.22
ViSAGe (FOV) Sphere360-Bench 210.87 2.90 1.51 0.71 1.49 73.45$\pm$1.42 74.89$\pm$1.71
ViSAGe (360) Sphere360-Bench 219.66 2.96 1.52 0.74 1.51 74.12$\pm$1.18 75.34$\pm$1.03
OmniAudio Sphere360-Bench 88.30 1.58 1.36 0.52 1.28 84.67$\pm$1.06 87.23$\pm$0.98
@@ 5.1 Experimental Setup
Baselines. Since our work pioneers in generating spatial audio from 360-degree videos, we construct the following baselines for comparison: (1) A cascaded system that integrates the state-of-the-art (video + text)-to-audio generation model MMAudio (Cheng et al., 2024a) with an audio spatialization component. Appendix E.4 includes details of the spatialization component. (2) Another cascaded system that integrates the classic video-to-audio generation model Diff-Foley (Luo et al., 2024) with the audio spatialization component. (3) ViSAGe (Anonymous, 2024), a recent model designed for generating spatial audio from perspective video inputs. We implement both ViSAGe with FOV video inputs and ViSAGe with panoramic videos for comparisons. ViSAGe is considered a competitive model as it specifically targets spatial audio generation. For all baselines except ViSAGe (360), we use the field-of-view of panoramic videos as video inputs, while for ViSAGe (360), we use the 360-degree video as input. Both baselines are trained on the Sphere360 dataset. We assess the performance of OmniAudio and baseline models on both the Sphere360-Bench and the YT360 test set. The YT360 test set is particularly valuable for evaluating the generalizability of our model due to its out-of-distribution characteristics. Note that audio spatialization is essentially the inverse process of calculating spatial angles: the spatialization is performed using the ground truth spatial angles. Therefore, “+AS” models are not evaluated using the three metrics related to evaluating spatial angles.
@@ 5.3 Case Study
In our qualitative analysis, we compare spectrograms of audio generated by OmniAudio and those produced by baselines, as illustrated in Figure 5. We make the following observations: (1) As demonstrated in case 1, FOV-based models struggle to generate spatial audio when the source object is out of view. In contrast, OmniAudio effectively generates FOA audio even when the object moves behind the camera, highlighting the importance of 360-degree video in accurately recreating the 3D sound environment required for FOA audio. (2) The spectrograms produced by OmniAudio consistently exhibit higher fidelity and alignment with video compared to those generated by baselines. In case 2, OmniAudio maintains a more consistent musical rhythm, as evidenced by the performance of the percussionist in the video. These enhancements are indicative of OmniAudio’s ability to better replicate the true acoustic scene, leading to more authentic and spatially precise audio outputs.
@@ 5.4 Ablation Study
Effect of Dual-branch Design. We investigate the effectiveness of the dual-branch design by comparing it with models that utilize only the input perspective video (w/ FOV), only the Equi-Angular Cubemap (a format of 360-degree video) (w/ EAC), and only equirectangular representations (w/ ERP). The results are presented in Table 4. We find that: (1) The dual-branch architecture outperforms single-branch models across all metrics, highlighting the benefit of integrating multiple representations for spatial audio generation. (2) While EAC outperforms ERP in single-branch settings, combining ERP and FOV captures complementary spatial information more effectively, likely because EAC’s division into six faces loses global context. (3) Panoramic videos significantly outperform FOV inputs in both spatial and non-spatial metrics, demonstrating the advantages of full-view 360-degree video for generating immersive spatial audio. By combining perspective video with equirectangular representations in the dual-branch design, the model

### v3

@@ Abstract
Traditional video-to-audio generation techniques primarily focus on perspective video and non-spatial audio, often missing the spatial cues necessary for accurately representing sound sources in 3D environments. To address this limitation, we introduce a novel task, 360V2SA, to generate spatial audio from 360-degree videos, specifically producing First-order Ambisonics (FOA) audio - a standard format for representing 3D spatial audio that captures sound directionality and enables realistic 3D audio reproduction. We first create Sphere360, a novel dataset tailored for this task that is curated from real-world data. We also design an efficient semi-automated pipeline for collecting and cleaning paired video-audio data. To generate spatial audio from 360-degree video, we propose a novel framework OmniAudio, which leverages self-supervised pre-training using both spatial audio data (in FOA format) and large-scale non-spatial data. Furthermore, OmniAudio features a dual-branch framework that utilizes both panoramic and perspective video inputs to capture comprehensive local and global information from 360-degree videos. Experimental results demonstrate that OmniAudio achieves state-of-the-art performance across both objective and subjective metrics on Sphere360. Code and datasets are available at github.com/liuhuadai/OmniAudio. The project website is available at OmniAudio-360V2SA.github.io.
video-to-audio generation, 360-degree video-to-spatial audio generation, spatial audio generation
{subcaptionblock}
0.49 Refer to caption {subcaptionblock}0.49 Refer to caption
Figure 1: Comparison of panoramic video and perspective video.
@@ 2 Related Work
Video-to-Audio Generation. This direction (Zhang et al., 2024; Wang et al., 2024a; Xu et al., 2024c; Tian et al., 2024) focuses on synthesizing audio that aligns seamlessly with the visual content of a video clip. Some approaches, such as SpecVQGAN (Iashin & Rahtu, 2021), FoleyGen (Mei et al., 2024), and V-AURA (Viertola et al., 2024), leverage autoregressive techniques to produce audio from silent video. Meanwhile, other methods (Luo et al., 2024; Cheng et al., 2024b; Xu et al., 2024b) adopt diffusion models (Song et al., 2020; Xing et al., 2024; Liu et al., 2024a) or flow matching (Lipman et al., 2022; Vyas et al., 2023; Wang et al., 2024b) generative models. Diff-Foley (Luo et al., 2024) employs an audio-visual contrastive feature and latent diffusion to predict spectrogram latent. Similarly, MovieGen Audio (Polyak et al., 2024) uses flow matching conditioned on multi-modal inputs, including videos and texts. Despite these advancements, a significant gap remains in the research focused on generating spatial audio from 360-degree video, an essential component for crafting genuinely immersive experiences. We introduce OmniAudio, which, to the best of our knowledge, represents the first approach in spatial audio generation from 360-degree video.
@@
Spatial Audio Generation. Existing spatial audio  [^1] Binaural audio captures sound using two microphones to simulate human ear positioning for a 3D effect when using headphones, while spatial audio involves a broader range of techniques that create 3D sound in dynamic environments, including with head tracking or multi-speaker setups. generation methods (Xu et al., 2021; Leng et al., 2022; Liu et al., 2023c; Gao & Grauman, 2019b; Kushwaha et al., 2024; Anonymous, 2024) predominantly focus on producing binaural (Sun et al., 2024; Yoshida et al., 2023; Garg et al., 2021) or First-order Ambisonics (FOA) audio (Kushwaha et al., 2024; Dagli et al., 2024) from fixed perspective inputs, such as mono audio, visual features, text, and source locations. For example, 2.5D Visual Sound (Gao & Grauman, 2019b) transforms monoaural audio into binaural sound using convolutional neural networks, while Diff-SAGe (Kushwaha et al., 2024) and ViSAGe (Anonymous, 2024) generate FOA audio by incorporating sound categories, source locations, and camera parameters. However, these approaches typically assume a static camera perspective, limiting their applicability in dynamic environments where the field of view can change in real-time. To overcome these limitations, we propose a framework for generating FOA audio from 360-degree video, which inherently supports dynamic and panoramic visual inputs. Our method leverages self-supervised flow matching pre-training and dual-branch video design to effectively synthesize spatial audio that adapts to the comprehensive and evolving visual context captured by 360-degree videos. This advancement enables more immersive and flexible audio-visual experiences in applications requiring real-time spatial audio adjustments.
@@
Flow Matching. Flow matching is used as the backbone for audio generation due to its superior generation performance. This framework (Lipman et al., 2022) trains a model to learn a vector field, enabling the generation of a desired probability path from noise to data. Unlike score-based models such as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), FM offers more stable and robust training and enhanced performance. This generative model has proven effective in audio generation, as seen in applications like Audiobox (Vyas et al., 2023) and FlashAudio (Liu et al., 2024b). Moreover, SpeechFlow (Liu et al., 2023a) has integrated self-supervised pre-training with FM to improve speech processing tasks, demonstrating its adaptability. In our work, we adapt FM for the novel task of spatial audio generation from 360-degree videos. By adopting a coarse-to-fine self-supervised pre-training strategy, our approach first learns general audio patterns from large-scale non-spatial audio data and then fine-tunes the model to capture spatial-specific characteristics. This enables the generation of high-quality First-order Ambisonics (FOA) audio that complements the immersive visual context of 360-degree videos.
@@ 4.1 Preliminaries
Conditional Flow Matching. Conditional flow matching is an effective generative technique applied across image (Esser et al., 2024; Lipman et al., 2022), audio (Iashin & Rahtu, 2021; Liu et al., 2024b), and video domains (Liu et al., 2024c). It employs a time-dependent velocity vector field parameterized by a neural network, denoted as $v_{\theta}(t,C,x)$, where $t$ represents time, $C$ is the conditioning input (e.g., video or text), and $x$ is a data point in $\mathbb{R}^{d}$.
@@ 4.2 Spatial Audio Representation
To address this gap and effectively encode FOA audio, we propose several modifications based on the Stable Audio framework: (1) We initialize our four-channel VAE with weights from a pre-trained stereo VAE to leverage existing non-spatial audio knowledge. (2) We eliminate the Mid-Side Short Time Fourier Transform (MS-STFT), which is specifically designed for stereo reconstruction, and adapt the system to transform the left/right components into FOA components W, X, Y, and Z, each with an loss weight of 1/4. For detailed information on VAE, please refer to Appendix D.
@@ 4.3 Dual-Branch Video Representation
We propose a dual-branch architecture to effectively model spatial-temporal dynamics in 360-degree videos. Given an input video sequence $V_{360}\in\mathbb{R}^{T\times H\times W\times C}$, which is in equirectangular projection format  [^3] The Equirectangular Projection (ERP) is a popular method for projecting 360-degree videos, commonly used by platforms like YouTube. It maps a spherical view onto a 2D image using latitude and longitude coordinates, with a 2:1 aspect ratio (360° horizontal, 180° vertical). with a height-to-width ratio of 1:2, we first pad it to a square format (1:1) and then resize it to match the input requirements of the image encoder. Here, $T$ is the number of frames, $H$ is the frame height, $W$ is the frame width, and $C$ is the number of color channels. We extract the perspective video  [^4] Perspective video denotes standard rectilinear projections with less than 120-degree horizontal field-of-view (FoV)., denoted as $V_{\text{fov}}\in\mathbb{R}^{T\times H^{\prime}\times W^{\prime}\times C}$, from the 360-degree video $V_{360}$ to capture local information. Our framework employs two complementary branches leveraging a frozen pre-trained MetaCLIP-Huge image encoder (Xu et al., 2024a):
@@ 5 Experiments
Table 2: Performance comparison between OmniAudio and the baselines on the Sphere360-Bench (in-distribution) and YT360 (out-of-distribution) test sets. We use objective metrics computing FD, KL divergence, $\Delta_{abs}\theta$, $\Delta_{abs}\phi$, and $\Delta_{\text{Angular}}$ between estimated DoA and ground truth, as well as subjective metrics including MOS for spatial audio quality (MOS-SQ) and video-audio alignment faithfulness (MOS-AF). We report the mean and standard deviation for MOS-SQ and MOS-AF. +AS denotes adding an audio spatialization component. For metrics with a downward arrow ($\downarrow$), lower values represent better performance, while for metrics with an upward arrow ($\uparrow$), higher values indicate better quality.
@@
In-distribution (Sphere360-Bench)
GT - - - - - - 88.41$\pm$0.79 90.12$\pm$1.08 -
Diff-Foley + AS 0.94B 331.05 3.56 - - - 69.87$\pm$0.84 71.12$\pm$1.36 2.40s
MMAudio + AS 1.03B 271.15 2.39 - - - 75.34$\pm$0.99 77.56$\pm$1.22 3.01s
ViSAGe (FoV) 0.36B 210.87 2.90 1.51 0.71 1.49 73.45$\pm$1.42 74.89$\pm$1.71 22.37s
ViSAGe (360) 0.36B 219.66 2.96 1.52 0.74 1.51 74.12$\pm$1.18 75.34$\pm$1.03 22.37s
OmniAudio 1.22B 88.30 1.58 1.36 0.52 1.28 84.67$\pm$1.06 87.23$\pm$0.98 0.92s
Out-of-distribution (YT360-Test)
GT - - - - - - 85.38$\pm$0.95 87.85$\pm$1.21 -
Diff-Foley + AS 0.94B 361.65 2.22 - - - 67.21$\pm$0.95 70.34$\pm$1.76 2.40s
MMAudio + AS 1.03B 190.40 1.71 - - - 73.25$\pm$1.05 76.77$\pm$1.23 3.01s
ViSAGe (FoV) 0.36B 199.09 1.86 2.21 0.88 1.99 71.82$\pm$1.98 72.17$\pm$1.47 22.37s
ViSAGe (360) 0.36B 225.52 1.95 2.18 0.86 1.98 72.45$\pm$1.64 72.96$\pm$1.39 22.37s
OmniAudio 1.22B 92.57 1.64 1.27 0.53 1.27 80.37$\pm$0.91 83.49$\pm$1.01 0.92s
@@ 5.1 Experimental Setup
Baselines. Since our work pioneers in generating spatial audio from 360-degree videos, we construct the following baselines for comparison: (1) A cascaded system that integrates the state-of-the-art (video + text)-to-audio generation model MMAudio (Cheng et al., 2024a) with an audio spatialization component. Appendix E.4 includes details of the spatialization component. (2) Another cascaded system that integrates the classic video-to-audio generation model Diff-Foley (Luo et al., 2024) with the audio spatialization component. (3) ViSAGe (Anonymous, 2024), a recent model designed for generating spatial audio from perspective video inputs. We implement both ViSAGe with FoV video  [^5] For all comparisons, we use the frontal-view large FoV (120°) as the default perspective video. This choice aligns with real-world conditions where front-facing viewpoints capture main sound sources. inputs and ViSAGe with panoramic videos for comparisons. ViSAGe is considered a competitive model as it specifically targets spatial audio generation. For all baselines except ViSAGe (360), we use the field-of-view of panoramic videos as video inputs, while for ViSAGe (360), we use the 360-degree video as input. Both baselines are trained on the Sphere360 dataset. We assess the performance of OmniAudio and baseline models on both the Sphere360-Bench and the YT360 test set. The YT360 test set is particularly valuable for evaluating the generalizability of our model due to its out-of-distribution characteristics. Note that audio spatialization is essentially the inverse process of calculating spatial angles: the spatialization is performed using the ground truth spatial angles. Therefore, “+AS” models are not evaluated using the three metrics related to evaluating spatial angles.
@@ 5.3 Case Study
In our qualitative analysis, we compare spectrograms of audio generated by OmniAudio and those produced by baselines, as illustrated in Figure 5. We make the following observations: (1) As demonstrated in case 1, MMAudio struggles to generate spatial audio when the source object is out of view. In contrast, OmniAudio effectively generates FOA audio even when the object moves behind the camera, highlighting the importance of 360-degree video in accurately recreating the 3D sound environment required for FOA audio. (2) The spectrograms produced by OmniAudio consistently exhibit higher fidelity and alignment with video compared to those generated by baselines. In case 2, OmniAudio maintains a more consistent musical rhythm, as evidenced by the performance of the percussionist in the video. These enhancements are indicative of OmniAudio’s ability to better replicate the true acoustic scene, leading to more authentic and spatially precise audio outputs.
@@ 5.4 Ablation Study
Effect of Dual-branch Design. We investigate the effectiveness of the dual-branch design by comparing it with models that utilize only the input perspective video (w/ Per), only the Equi-Angular Cubemap (a format of 360-degree video) (w/ EAC), and only equirectangular representations (w/ ERP). The results are presented in Table 4. We find that: (1) The dual-branch architecture outperforms single-branch models across all metrics, highlighting the benefit of integrating multiple representations for spatial audio generation. (2) While EAC outperforms ERP in single-branch settings, combining ERP and Per captures complementary spatial information more effectively, likely because EAC’s division into six faces loses global context. (3) Panoramic videos significantly outperform perspective inputs in both spatial and non-spatial metrics, demonstrating the advantages of full-view 360-degree video for generating immersive spatial audio. By combining perspective video with equirectangular representations in the dual-branch design, the model